# Optomechanical Systems with Linear and Quadratic Position Couplings: Dynamics and Optimal Estimation

### v1

@@ IV.2 Real-world system
with $r=0.42$, from which we can calculate the expansion expressed in Eq. (2). This setup guarantees a relatively high magnitude of the quadratic optomechanical coupling $g_{2}\approx 2\pi\times 24\text{ kHz}\cdot\text{nm}^{-2}$ obtained by placing the membrane either in a node, or in an antinode of the magnetic field, when $g_{1}\approx 0$. Furthermore, in Refs.  [36, 37] it is shown how, by tilting the membrane of a small angle, we can induce an interaction between different transverse modes of the electromagnetic field, leading to an increased coupling constant $g_{2}\approx 2\pi\times 4.46\text{ MHz}\cdot\text{nm}^{-2}$. We consider the system to be at room temperature $T=300K$, and we consider the initial state of the mechanical system, (15) to be a thermal state with $n_{\text{th}}\approx 10^{5}$, Eq. (32). These values are reported in Table 1.
@@
Using these values, we found that the first peak of the QFI is $\approx 5.5\times 10^{15}$ for an initial mixed state $\hat{\rho}_{s}$ with $s=0.2$. We analyze how the choice of the parameter $\phi$ influences the CFI. While the behavior shown in Fig. 8 is reflected even with the real-world values, we see that a different choice of $\phi$ can lead the CFI to saturate the QFI. In Fig. 9 we plot at $t=2\pi(\frac{1}{\Delta}+\frac{1}{\delta})$, the CFI as a function of the phase $\phi.$
Figure 9: QFI (red dashed line) and CFI (black continuous line) for the real-world system, calculated at $t=2\pi(\frac{1}{\Delta}+\frac{1}{\delta})$ as a function of the quadrature phase $\phi$.
Finally, we investigate the influence of temperature on the QFI. In Fig. 10 we plot the QFI as a function of the temperature for the different peaks at $t=2\pi N/\Delta,$ with $N=1,10,$ and $100$
@@
Figure 10: Quantum Fisher Information as a function of the system temperature plotted at $t=2\pi N/\Delta$, with $N=1,10,$ and $100$. Depending on the temperature the largest peak of the QFI is located at different periods of oscillation.
@@
When the temperature is close to zero the state of the mechanical system is substantially the ground state of the mechanical oscillator. This means that the optical system is interacting with a very weak field and consequently the information about the optomechanical coupling is low. At the opposite, when the temperature is large, the mechanical system is close to the maximally mixed state, which is persistent, regardless of the value of $g_{2}.$ Even in this case, the information retrieval about $g_{2}$ is small. In between, there is a temperature where the QFI is at its maximum.
@@
From Fig. 10 we see that the largest peak of the QFI is located at a different period of oscillation that depends on the temperature of the mechanical system. This behavior is opposite for cold and hot temperatures. For low $T,$ although the mechanical system is approximately in the ground state, the information grows at each cycle, as the information available to the optical system cumulates. For high $T,$ the mechanical bath tends to destroy all the available information and the peaks of the QFI decline at each cycle.
@@ V Conclusion and perspectives
In this work we studied the dynamics of an optomechanical system with quadratic coupling in the position operator. We found that the each photon number brings a separate contribution to the global Hamiltonian, thus allowing us to isolate these contributors and solve the temporal evolution of the density operator. Under the assumption that the optical and mechanical quantum states are factorized at initial time, and considering the mechanical system in a thermal state, we were able to derive the full description of the optical density matrix.
We studied the evolution on the Bloch sphere of a simple system with the initial optical state being a superposition of zero and one-photon states. The optical system has a purity that oscillates with a period frequency given by the beat frequency between mechanical system associated with different photons number.
Hence, we analyzed the sensitivity of the optical field to changes of the quadratic optomechanical coupling by calculating the quantum Fisher information on the reduced density matrix. The effect of the quadratic optomechanical coupling becomes evident when the system regains its original purity, as the off-diagonal elements of the optical density matrix store the information about the coupling constant. We studied the classical Fisher information associated with the measurement of the quadrature operator, finding that it is able to saturate the quantum Fisher information when the phase of the quadrature and the phase aligns with the phase of the off-diagonal terms.
Furthermore, we studied the dependence of the quantum Fisher information to the temperature of the mechanical system, showing that it reaches its maximum at an intermediate temperature that allows the state of the mechanical system to be highly populated while being far from the maximally mixed state.

### v2

@@ IV.2 Interdependent couplings
For the results presented in this subsection, we use the numerical values of the system parameters reported in [37]. The setup consists of a suspended SiN membrane oscillating at a frequency of $\Omega=2\pi\times 134\,\text{kHz}$, positioned at the center of a Fabry-Pérot optical cavity of length $L=67\text{mm}$, and driven by an external laser with wavelength $\lambda=1064\,\text{nm}$. This configuration yields an effective harmonic oscillator mass of approximately $m\approx 50\,\text{pg}$. The membrane has a reflectivity of $r=0.42$, and the system operates at a temperature of $300\,\text{K}$. Under these conditions, the initial thermal state of the mechanical oscillator described in Eq. (18) corresponds to an average phonon number of $n_{\text{th}}\approx 4.665\times 10^{7}$. This setup yields a relatively large quadratic optomechanical coupling of $g_{2}\approx 1.445\times 10^{23}\,\text{Hz}\cdot\text{m}^{-2}$, as obtained from Eq. (69) when the membrane is positioned at either a node or an antinode of the optical field, i.e., $x_{0}=n\lambda/2$, with $n=400$. The remaining parameters are determined using Eqs. (67) and (68) and are summarized in Table 1. Furthermore, Refs. [33, 14] show that tilting the membrane by a small angle can induce coupling between different transverse modes of the electromagnetic field, thereby enhancing the quadratic coupling constant to $g_{2}\approx 2.802\times 10^{25}\,\text{Hz}\cdot\text{m}^{-2}$.
IV.2.1 Estimates of the quadratic optomechanical coupling
We aim to estimate the value of the optomechanical coupling constant, denoted by $g$, in an experimental context. To treat this as an unknown parameter, we assume that the phenomenological model given in Eq. (64) provides a good, though approximate, description of the coupling. To investigate this scenario, we use the parameter values listed in Table 1. Note that throughout this subsection, we exclusively consider the mixed initial state corresponding to case (ii) in Section IV.1.
We compute and plot the QFI over a range of values of the quadratic coupling parameter $g$, centered around the phenomenological reference value $g_{2}$. For convenience, we introduce the notation $\Omega_{n,g}=\sqrt{\Omega^{2}+\frac{2\hbar g_{2}n}{m}}$ to emphasize the dependence of $\Omega_{n}$ on $g$ in Eq. (7). Fig. 8 presents three data series, each showing the QFI evaluated for different values of $g$ (on the $x$-axis) and at the reference peaking time associated with three specific choices of $g$, namely $0.95g_{2}, g=g$$g_{2}$, and $g=1.05g_{2}$. The corresponding times are given by $t=T_{\Omega}=\frac{2\pi}{\Omega_{1,g}-\Omega_{0,g}}$.
The results show that, for each selected system time, the QFI exhibits a distinct sharp peak precisely centered at the corresponding $g$ value. This indicates a strong correlation between the system’s information-theoretic properties and the characteristic reference time $T_{\Omega}$ in a simple two-level system composed of zero- and one-photon Fock states. Furthermore, by comparing with Fig. 7, we observe that as more Fock states participate in the system dynamics, the periodicity of the QFI decreases, and the peaks no longer appear at integer multiples of $T_{\Omega}$.
Figure 8: Quantum Fisher information of an experimental system as a function of the phenomenological values of $g_{2}$, evaluated at three different time points: $t=T_{g,\Omega}=\frac{2\pi}{\Omega_{1,g}-\Omega_{0,g}}$ for $g=0.95g_{2}$ (blue dashed line), $g=g_{2}$ (red solid line), and $g=1.05g_{2}$ (green dotted line). The range of the unknown coupling values $g$ considered spans from $0.9g_{2}$ to $1.1g_{2}$. All other system parameters are kept fixed at the values listed in Table 1.
@@ IV.2.1 Estimates of the quadratic optomechanical coupling
Using these values, we found that the first peak of the QFI is $\approx 5.5\times 10^{15}$ for an initial mixed state $\hat{\rho}_{s}$ with $s=0.2$. We analyze how the choice of the parameter $\phi$ influences the CFI. While the behavior shown in Fig. 6 is reflected even with the experimental values, we see that a different choice of $\phi$ can lead the CFI to saturate the QFI. In Fig. 9 we plot at $t=2T_{\Omega}$, the CFI as a function of the phase $\phi$.
Figure 9: Quantum Fisher information (red dashed line) and classical Fisher information (black continuous line) for the experimental system, calculated at $t=2T_{\Omega}$ as a function of the phase $\phi$. System parameters take on the values in Table 1.
@@
Figure 10: Quantum Fisher information as a function of the system temperature plotted at $t=N\times T_{\Omega}$, with $N=10$, $100$, and $1000$. Depending on the temperature the largest peak of the QFI is located at different periods of oscillation. The other system parameters are given in Table 1.
@@
Finally, we investigate the influence of temperature on the QFI. In Fig. 10 we plot the QFI as a function of the temperature for the different peaks at $t=N\times T_{\Omega}$ with $N=10$, $10^{3}$, and $10^{5}$. When the temperature is close to zero the state of the mechanical system is substantially the ground state of the mechanical oscillator. This means that the optical system is interacting with a very weak field and consequently the information about the optomechanical coupling is low. At the opposite, when the temperature is large, the mechanical system is close to the maximally mixed state, which is persistent, regardless of the value of $g$. Even in this case, the information retrieval about $g$ is small. In between, there is a temperature where the QFI is at its maximum.
@@
From Fig. 10 we see that the largest peak of the QFI is located at a different oscillation period that depends on the temperature of the mechanical system. This behavior is opposite for cold and hot temperatures. For low $T$ although the mechanical system is approximately in the ground state, the information grows at each cycle, as the information available to the optical system cumulates. For high temperature $T,$ the mechanical bath tends to destroy all the available information and the peaks of the QFI decline at each cycle.
@@ V Conclusion and perspectives
In this work, we have investigated the dynamics of an optomechanical system with linear and quadratic couplings. We have found that each photon number contributes independently to the total Hamiltonian, which allows us to isolate these contributions and solve the temporal evolution of the density operator. Assuming that the optical and mechanical quantum states are initially factorized, we have derived a complete description of the optical field’s density operator using the formalism of two-photon coherent states. Furthermore, by considering the mechanical system to be initially in a thermal state, we obtained a solution that can be evaluated numerically.
We have investigated the dynamics of the system for two types of initial optical field states: a coherent state, and a state restricted to the subspace spanned by the zero- and one-photon Fock states. To understand the influence of the linear and quadratic optomechanical couplings on the system’s evolution, we first considered a dimensionless model in which the two couplings are treated as independent.
The evolution of the optical field has been analyzed using the Wigner phase-space distribution and, for the two-level subspace, the Bloch sphere representation. The Wigner distribution reveals a rich dynamics in which the photon number is preserved, while intricate interference fringes develop. In the two-level case, through a series of comprehensive analysis of the system dynamics that considers different combinations of $g_{1}$ and $g_{2}$, we find strong approximate periodicity in the system when one of the couplings is suppressed. As manifested in the experimentally relevant regime, where $g_{1}\gg g_{2}$, the periodicity are attributed to the photon number-dependent frequencies of the mechanical oscillator.
To quantify the sensitivity of the optical field to variations in the quadratic coupling, we have calculated the quantum Fisher information for both initial conditions of the field’s quantum state. The effect of the quadratic optomechanical coupling is particularly evident when the system regains its original purity, as the off-diagonal elements of the optical density matrix encode information about the coupling constant. Additionally, we have evaluated for the two-level case the classical Fisher information associated with balanced homodyne photodetection and found that it is capable of saturating the quantum Fisher information.
Finally, we have considered a phenomenological model for the cavity frequency, which results in interdependent linear and quadratic couplings. Using the experimental values from Eq. [37], we have investigated the estimation of the quadratic coupling around the value predicted by the phenomenological model. Consistent with the findings from the dimensionless model, we have also demonstrated that in a balanced homodyne photodetection setup, adjusting the phase of the coherent local oscillator allows the classical Fisher information to saturate the quantum Fisher information. Furthermore, we have studied the dependence of the quantum Fisher information to the temperature of the mechanical system, showing that it reaches its maximum at an intermediate temperature that allows the state of the mechanical system to be highly populated while being far from the maximally mixed state.